Executive salaries are charged to three operating divisions on the basis of capital ivestment in the three divisions. I the investment is 315.8 milion in the Northern Division, 59.4 milion in the Eastem Divieion, and S12.6 million in the Western Division, how should executive salaries of 5642,600 be allocated to the three divisions? visions alaries orthern astem astem Round to the nearest dollar as needad)
Answer #1

Investment in Northern division = $15.8 m

Eastern division = $9.4 m

Western division = $12.6 m

Total investment = (15.8+9.4+ 12.6) = 37.8m

Salary allocated to Northern division = Total salary* Investment in Northern division/ Total investment

= 642600*15.8/37.8

=268600

Salary allocated to Eastern division = Total salary* Investment in Eastern division/ Total investment

= 642600*9.4/37.8

=159800

Salary allocated to Western division = Total salary* Investment in Western division/ Total investment

= 642600*12.6/37.8

=214200